Sue works a desk job, commutes to work, parks close to the building and watches television with her family in the evening. At 165 pounds, she expends about 2,000 calories a day. Her average calorie intake is 2,150 calories per day. If one pound of body fat is equal to approximately 3500 extra calories, how many days will it take for Sue to gain a pound of body fat if she maintains 2150 calories per day?

*** This is all the information I was given***
Answer #1

Sue's calorie intake = 2150Cal

her energy expenditure = 2000 Cal

Calorie storage for her = 2150-2000 =150 Calories

1 pound of body fat = 3500 Calories

how many days she requires to store 3500 Calorie?

   3500/150 =23.33 or 24 days